Suppose the difference between matter and antimatter in the early universe were 1 part in 108 instead of 1 part in 109.
(a.) Evaluate the temperature at which deuterium begins to form.
(b.) At what age does this occur?
(c.) Evaluate the temperature and the corresponding time of radiation decoupling (when hydrogen atoms form).
Answer #1

25 f=-710-8-0.42e (a) + 2 gives kT kT kT 2.22 x10° eV 1.0×109 K T=_= 25(8.617×10-5eV/K) 25k 1.5×1010 s1/2 . K 1.5×1010 s1/2K

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions